# Observational Constraints on Dark Energy Models with $\Lambda$ as an Equilibrium Point

### v1

@@ V.2 $SN+CC+BAO$ data
We continue with the second family of models, which follows from the logarithmic parametrization. These models do not fit the cosmological data in the same way. The corresponding values for $\chi_{\min}^{2}$ are $\chi_{\min}^{2}\left(w_{+}^{IV}\right)=1435.1$, $\chi_{\min}^{2}\left(w_{-}^{IV}\right)=1434.8$, $\chi_{\min}^{2}\left(w^{V}\right)=1435.4$ and $\chi_{\min}^{2}\left(w^{VI}\right)=1434.9$. Thus, for the models $w_{\pm}^{IV}$, we calculate $\Delta\left(AIC\right)<2$, indicating that these models are statistically equivalent to the $\Lambda$CDM model. The results are summarized in Table 3.
When BAO data are introduced, models $w_{\pm}^{IV}\left(z\right)$ and $w^{VI}\left(z\right)$ lead to a smaller value of $\Delta\left(AIC\right)$ in comparison to the $\Lambda$CDM model. However, this is not the case for model $w^{V}\left(z\right)$. We note that the parameters of model $w^{VI}\left(z\right)$ cannot be constrained by the data analysis. Models $w^{V}\left(z\right)$ and $w^{VI}\left(z\right)$ are ruled out by these cosmological data.
@@
Table 3: Constraints for the $SN+CC+BAO$ data.
$\Lambda$CDM $69.1_{-1.5}^{+1.5}$ $0.26_{-0.012}^{+0.012}$ $\nexists$ $\nexists$ $1435.4$ 00
$w_{+}^{I}\left(z\right)$ $69.0_{-1.8}^{+1.4}$ $0.26_{-0.021}^{+0.024}$ $0.04_{-0.37}^{+0.37}$ $\nexists$ $1436.4$ $3.0$
$w_{-}^{I}\left(z\right)$ $68.7_{-1.6}^{+1.6}$ $0.25_{-0.020}^{+0.029}$ $-0.35_{-0.40}^{+0.47}$ $\nexists$ $1435.6$ $2.2$
$w^{II}\left(z\right)$ $68.8_{-1.6}^{+1.2}$ $0.25_{-0.019}^{+0.023}$ $0.17_{-0.30}^{+0.38}$ $\nexists$ $1435.7$ $2.1$
$w^{III}\left(z\right)$ $68.7_{-1.3}^{+1.1}$ $0.26_{-0.013}^{+0.013}$ $-16.6_{-7.3}^{+7.3}$ $2.0_{-1.0}^{+2.7}$ $1433.6$ $2.2$
$w_{+}^{IV}\left(z\right)$ $68.7_{-1.4}^{+1.4}$ $0.26_{-0.012}^{+0.012}$ $>8.6$ $\nexists$ $1435.1$ $1.7$
$w_{-}^{IV}\left(z\right)$ $68.7_{-1.2}^{+1.3}$ $0.27_{-0.012}^{+0.012}$ $9.9_{-8.8}^{+7.1}$ $\nexists$ $1434.8$ $1.4$
$w^{V}\left(z\right)$ $68.4_{+1.7}^{+1.7}$ $0.26_{-0.009}^{+0.009}$ $>6.47$ $7.3_{-5.7}^{+6.8}$ $1435.4$ $4$
$w^{VI}\left(z\right)$ $68.9_{+1.5}^{+1.2}$ $0.26_{-0.013}^{+0.011}$- -- $1434.9$ $3.5$
@@
Figure 4: Contour plots for $w^{I}_{+}$ $w^{I}_{-}$
@@ VI Conclusions
On the other hand, for the second family of models, which follows from the logarithmic parametrization, we found that $w_{\pm}^{IV}\left(z\right)$ models are statistically equivalent to $\Lambda$CDM with $\Delta\left(AIC\right)<2$. However, models $w^{V}\left(z\right)$ and $w^{VI}\left(z\right)~{}$were strongly disfavored by the data. The parameters of model $w^{VI}\left(z\right)~{}$could not be constrained, which means that there exists a high degree of correlation in the parameter space.
Table 4: Constraints for the $SN+CC+BAO$ data of modified CPL models.
Model $\mathbf{H}_{0}$ $\mathbf{\Omega}_{m0}$ $\mathbf{w}_{0}$ $\mathbf{w}_{a}$ $\mathbf{w}_{b}$ $\chi_{\min}^{2}$ $\mathbf{AIC-AIC}_{CPL}$ $\mathbf{AIC-AIC}_{\Lambda}$
CPL $68.4_{-1.4}^{+1.1}$ $0.26_{-0.015}^{+0.021}$ $-0.90_{-0.056}^{+0.056}$ $-0.21_{-0.55}^{+0.22}$ $\nexists$ $1433.3$ 00 $1.9$
$\hat{w}_{+}^{I}\left(z\right)$ $67.9_{-1.6}^{+1.6}$ $0.25_{-0.009}^{+0.009}$ $-0.90_{-0.053}^{+0.053}$ $0.01_{-0.21}^{+0.19}$ $\nexists$ $1433.9$ $0.6$ $2.5$
$\hat{w}_{-}^{I}\left(z\right)$ $67.8_{-1.6}^{+1.6}$ $0.25_{-0.009}^{+0.009}$ $-0.90_{-0.053}^{+0.053}$ $0.02_{-0.2}^{+0.2}$ $\nexists$ $1433.9$ $0.6$ $2.5$
$\hat{w}^{II}\left(z\right)$ $68.5_{-1.2}^{+1.2}$ $0.26_{-0.016}^{+0.020}$ $-0.89_{-0.066}^{+0.050}$ $-0.40_{-0.34}^{+0.57}$ $\nexists$ $1433.1$ $-0.2$ $1.7$
$\hat{w}^{III}\left(z\right)$ $68.4_{-1.2}^{+1.2}$ $0.25_{-0.065}^{+0.034}$ $-0.94_{-0.065}^{+0.034}$ $<-16.0$ $0.7_{-4.2}^{+4.2}$ $1433.1$ $1.8$ $3.7$
$\hat{w}_{+}^{IV}\left(z\right)$ $68.4_{-1.1}^{+1.1}$ $0.25_{-0.014}^{+0.014}$ $-0.93_{-0.039}^{+0.046}$ $11.2_{-6.3}^{+5.0}$ $\nexists$ $1432.9$ $-0.4$ $1.5$
$\hat{w}_{-}^{IV}\left(z\right)$ $68.4_{-1.2}^{+1.2}$ $0.26_{-0.014}^{+0.014}$ $-0.93_{-0.038}^{+0.038}$ $10.1_{-5.5}^{+5.5}$ $\nexists$ $1433.0$ $-0.3$ $1.6$
$\hat{w}^{V}\left(z\right)$ $68.0_{-1.7}^{+1.7}$ $0.25_{-0.014}^{+0.020}$ $-0.92_{-0.055}^{+0.042}$ $8.0_{-5.7}^{+3.4}$ $>5.2$ $1433.7$ $2.4$ $4.3$
$\hat{w}^{VI}\left(z\right)$ $68.5_{-1.4}^{+1.2}$ $0.25_{-0.014}^{+0.014}$ $-0.92_{-0.052}^{+0.041}$ $>10.3$- $1433.3$ $2$ $3.9$
@@
We observe that most models fit the observational data with a $\chi_{\min}^{2}$ value smaller than that of CPL and $\Lambda$CDM. However, due to the different number of degrees of freedom, the application of the AIC indicates that, in comparison to CPL, the models $\hat{w}^{A}\left(z\right)$ are statistically indistinguishable. The best fit is provided by $\hat{w}_{-}^{IV}\left(z\right)$, which is known as the $n$CPL model.
Although these models have been introduced to describe the asymptotic behavior of $w\left(z\right)$ near the equilibrium point at the present time, our analysis shows that these models can be used for observational constraints in the late universe. However, for CMB data, where $z\gg 1$, we show that the free parameters are degenerate, and this approximation may not be valid. Nevertheless, by reconstructing the asymptotic behavior of $w\left(z\right)$ at different epochs, we can further constrain the physical law that $w\left(z\right)$ satisfies. This will provide with important insights into the nature of dark energy.
@@ Acknowledgements.
AP thanks the support of VRIDT through Resolución VRIDT No. 096/2022 and Resolución VRIDT No. 098/2022. Part of this study was supported by FONDECYT 1240514. AP thanks Dr. F. Anagnostopoulos for a fruitful discussion.

### v2

@@ V.2 DESI  DR1: $SN+CC+BAO_{1}$ data
We continue with the second family of models, which follows from the logarithmic parametrization. These models do not fit the cosmological data in the same way. The corresponding values for $\chi_{\min}^{2}$ are $\chi_{\min}^{2}\left(w_{+}^{IV}\right)=1435.3$, $\chi_{\min}^{2}\left(w_{-}^{IV}\right)=1434.8$, $\chi_{\min}^{2}\left(w^{V}\right)=1435.3$ and $\chi_{\min}^{2}\left(w^{VI}\right)=1435.0$. Thus, for the models $w_{\pm}^{IV}$, we calculate $\Delta\left(AIC\right)\leq 2$, indicating that these models are statistically equivalent to the $\Lambda$CDM model. The results are summarized in Table 3.
When BAO1 data are introduced, models $w_{\pm}^{IV}\left(z\right)$ and $w^{VI}\left(z\right)$ lead to a smaller value of $\Delta\left(AIC\right)$ in comparison to the $\Lambda$CDM model. However, this is not the case for model $w^{V}\left(z\right)$. We note that the parameters of model $w^{VI}\left(z\right)$ cannot be constrained by the data analysis. Models $w^{V}\left(z\right)$ and $w^{VI}\left(z\right)$ are ruled out by these cosmological data.
Table 3: Constraints for the $SN+CC+BAO_{1}$ data.
$\Lambda$CDM $68.8_{-1.7}^{+1.2}$ $0.311_{-0.012}^{+0.012}$ $\nexists$ $\nexists$ $1435.2$ 00
$w_{+}^{I}\left(z\right)$ $68.7_{-1.6}^{+1.2}$ $0.303_{-0.018}^{+0.024}$ $0.09_{-0.27}^{+0.39}$ $\nexists$ $1435.8$ $2.6$
$w_{-}^{I}\left(z\right)$ $68.8_{-1.6}^{+1.2}$ $0.293_{-0.019}^{+0.028}$ $-0.34_{-0.36}^{+0.47}$ $\nexists$ $1435.4$ $2.2$
$w^{II}\left(z\right)$ $68.8_{-1.6}^{+1.2}$ $0.298_{-0.017}^{+0.029}$ $0.20_{-0.37}^{+0.37}$ $\nexists$ $1435.8$ $2.6$
$w^{III}\left(z\right)$ $68.7_{-1.2}^{+1.2}$ $0.307_{-0.012}^{+0.012}$ $-17.1_{-9.0}^{+6.8}$ $>1.57$ $1433.5$ $2.3$
$w_{+}^{IV}\left(z\right)$ $68.6_{-1.5}^{+1.3}$ $0.310_{-0.012}^{+0.012}$  $\,11.3_{-4.8}^{+4.8}$ $\nexists$ $1435.3$ $2$
$w_{-}^{IV}\left(z\right)$ $68.6_{-1.4}^{+1.4}$ $0.313_{-0.012}^{+0.012}$ $9.8_{-7.7}^{+6.6}$ $\nexists$ $1434.8$ $1.6$
$w^{V}\left(z\right)$ $68.5_{-1.6}^{+1.6}$ $0.310_{-0.014}^{+0.014}$ $>5.81$ $>3.90$ $1435.3$ $4.1$
$w^{VI}\left(z\right)$ $69.1_{-1.5}^{+1.2}$ $0.312_{-0.012}^{+0.012}$- -- $1435.0$ $3.8$
V.3 DESI  DR2:${}SN+CC+BAO_{2}$ data
We introduce the BAO2 data as given by the DESI DR2 and we perform the same analysis. We show that for these models there is not any difference between the BAO1 and BAO2 data in terms of the $\chi_{\min}^{2}$ or the $\Delta\left(AIC\right)$. $\Lambda$CDM fits the data in a better way with the rest of the models. However, there is small statistical preference on the $\Lambda$CDM in comparison with the rest of the parametric dark energy models.
Table 4: Constraints for the $SN+CC+BAO_{2}$ data.
$\Lambda$CDM $68.9_{-1.5}^{+1.2}$ $0.311_{-0.012}^{+0.012}$ $\nexists$ $\nexists$ $1435.1$ 00
$w_{+}^{I}\left(z\right)$ $68.8_{-1.5}^{+1.2}$ $0.303_{-0.018}^{+0.023}$ $0.12_{-0.28}^{+0.35}$ $\nexists$ $1435.7$ $2.6$
$w_{-}^{I}\left(z\right)$ $68.8_{-1.6}^{+1.2}$ $0.293_{-0.020}^{+0.031}$ $-0.33_{-0.42}^{+0.51}$ $\nexists$ $1435.8$ $2.7$
$w^{II}\left(z\right)$ $68.8_{-1.6}^{+1.3}$ $0.296_{-0.019}^{+0.028}$ $0.25_{-0.35}^{+0.35}$ $\nexists$ $1435.9$ $2.8$
$w^{III}\left(z\right)$ $68.6_{-1.4}^{+1.0}$ $0.308_{-0.012}^{+0.012}$ $-17.1_{-8.5}^{+7.2}$ $2.6_{-1.5}^{+2.9}$ $1433.4$ $2.3$
$w_{+}^{IV}\left(z\right)$ $068.6_{-1.5}^{+1.3}$ 0 $0.310_{-0.012}^{+0.012}$ $\,>7.61$ $\nexists$ $1435.3$ $2.1$
$w_{-}^{IV}\left(z\right)$ $68.8_{-1.4}^{1.4}$ $0.313_{-0.012}^{+0.012}$ $9.1_{-7.3}^{+3.1}$ $\nexists$ $1434.9$ $1.8$
$w^{V}\left(z\right)$ $68.5_{-1.7}^{+1.7}$ $0.312_{-0.014}^{+0.014}$ $>6.65$ $>4.39$ $1435.4$ $4.3$
$w^{VI}\left(z\right)$ $68.9_{-1.4}^{+1.2}$ $0.311_{-0.012}^{+0.012}$ $14.5_{-8.5}^{+8.5}$ $>-2.84$ $1434.7$ $3.6$
@@ V.3 DESI  DR2:${}SN+CC+BAO_{2}$ data
Figure 4: Contour plots for $w_{+}^{I}$ $w_{-}^{I}$
@@ VI Conclusions
On the other hand, for the second family of models, which follows from the logarithmic parametrization, we found that $w_{\pm}^{IV}\left(z\right)$ models are statistically equivalent to $\Lambda$CDM with $\Delta\left(AIC\right)\leq 2$. However, models $w^{V}\left(z\right)$ and $w^{VI}\left(z\right)~{}$were strongly disfavored by the data. The parameters of model $w^{V}\left(z\right)~{}$could not be constrained, which means that there exists a high degree of correlation in the parameter space.
Table 5: Constraints for the $SN+CC+BAO_{2}$ data of modified CPL models.
Model $\mathbf{H}_{0}$ $\mathbf{\Omega}_{m0}$ $\mathbf{w}_{0}$ $\mathbf{w}_{a}$ $\mathbf{w}_{b}$ $\chi_{\min}^{2}$ $\mathbf{AIC-AIC}_{CPL}$ $\mathbf{AIC}-\mathbf{AIC}_{\Lambda}$
CPL $68.4_{-1.4}^{+1.1}$ $0.308_{-0.015}^{+0.021}$ $-0.893_{-0.056}^{+0.056}$ $-0.33_{-0.48}^{+0.43}$ $\nexists$ $1433.3$ 00 $2.2$
$\hat{w}_{+}^{I}\left(z\right)$ $68.2_{-1.4}^{+1.2}$ $0.301_{-0.012}^{+0.026}$ $-0.894_{-0.057}^{+0.057}$ $-0.26_{-0.74}^{+0.26}$ $\nexists$ $1433.4$ $0.1$ $2.3$
$\hat{w}_{-}^{I}\left(z\right)$ $68.3_{-1.4}^{+1.2}$ $0.303_{-0.016}^{+0.028}$ $-0.906_{-0.060}^{+0.060}$ $0.17_{-0.36}^{+0.65}$ $\nexists$ $1433.6$ $0.3$ $2.5$
$\hat{w}^{II}\left(z\right)$ $68.4_{-1.4}^{+1.1}$ $0.311_{-0.017}^{+0.021}$ $-0.885_{-0.061}^{+0.061}$ $-0.43_{-0.42}^{+0.56}$ $\nexists$ $1433.5$ $0.2$ $2.4$
$\hat{w}^{III}\left(z\right)$ $68.3_{-1.1}^{+1.1}$ $0.300_{-0.013}^{+0.013}$ $-0.993_{-0.056}^{+0.042}$ $<16.9$ $0.98_{-4.3}^{+4.8}$ $1433.2$ $1.9$ $4.1$
$\hat{w}_{+}^{IV}\left(z\right)$ $68.3_{-1.2}^{+1.2}$ $0.295_{-0.013}^{+0.017}$ $-0.914_{-0.053}^{+0.044}$ $>7.4$ $\nexists$ $1433.1$ $-0.2$ $2.0$
$\hat{w}_{-}^{IV}\left(z\right)$ $68.4_{-1.2}^{+1.2}$ $0.302_{-0.015}^{+0.015}$ $-0.922_{-0.051}^{+0.040}$ $>6.7$ $\nexists$ $1433.1$ $-0.2$ $2.0$
$\hat{w}^{V}\left(z\right)$ $68.0_{-1.6}^{+1.6}$ $0.298_{-0.015}^{+0.018}$ $-0.916_{-0.053}^{+0.044}$ $>5.6$ $7.4_{-6.1}^{+3.0}$ $1433.8$ $2.5$ $4.7$
$\hat{w}^{VI}\left(z\right)$ $68.4_{-1.4}^{+1.1}$ $0.299_{-0.050}^{+0.042}$ $-0.920_{-0.050}^{+0.042}$ $16_{-11}^{+14}$- $1433.2$ $1.9$ $4.1$
@@
We observe that most models fit the observational data with a $\chi_{\min}^{2}$ value smaller than that of CPL and $\Lambda$CDM. However, due to the different number of degrees of freedom, the application of the AIC indicates that, in comparison to CPL, the models $\hat{w}^{A}\left(z\right)$ are statistically indistinguishable. The best fit is provided by $\hat{w}_{-}^{IV}\left(z\right)$, which is known as the $n$CPL model. Finally, if when we compare the results of Table 4 and 5, we conclude that the DESI DR2 BAO data support models with $w_{0}>-1$.
Although these models have been introduced to describe the asymptotic behavior of $w\left(z\right)$ near the equilibrium point at the present time, our analysis shows that these models can be used for observational constraints in the late universe. However, for CMB data, where $z\gg 1$, we show that the free parameters are degenerate, and this approximation may not be valid.
This consideration is different from the Taylor expansion of the $w\left(a\right)$ around the present time. Indeed, the CPL can be seen as the first-order approximation of a nonlinear $w\left(z\right)$ in the Taylor expansion. The introduction of additional terms in the Taylor expansion is done with the charge of additional free parameters. This can lead to new problems as discussed recently in bb11 . However, in our approach we can reconstructing the asymptotic behavior of $w\left(z\right)$ at different epochs. This can lead to further constrain to the physical law that $w\left(z\right)$ satisfies. In turn, can provide important insights into the nature of dark energy. From this work it follows that the logarithmic parametrization leads to parametric dark energy models which fit the datta in a better way. Moreover, the existence of oscillations, i.e. model $\hat{w}^{VI}\left(z\right)$, lead to smaller values of the $\chi_{\min}^{2}$, with the cost of a greater value for the $AIC$ due to the larger number of free parmameters. We recall that the Chiral multi-scalar field theory ch1 ; ch2 ; ch3 ; ch4 , known as warm inflation ch5 , provide such oscillationg behaviours.
@@ Acknowledgements.
AP thanks the support of VRIDT through Resolución VRIDT No. 096/2022 and Resolución VRIDT No. 098/2022. Part of this study was supported by FONDECYT 1240514 Etapa 2025. AP thanks Dr. F. Anagnostopoulos for a fruitful discussion.